# Antonio Ramírez-de-Arellano

### v2

@@ 3.5 Discussion of old ingredients
Singleton (Theorem 3.5) $=$ 1 1 1
$\mathrm{NLinFIN}$ (Proposition 3.7) $=$ 2 2 *
$\mathrm{NCombFIN}$ (Proposition 3.9) $=$ 3 3 *
Finite sets [4] $\subseteq$ 1 * *
[4] $\subseteq$ * * 1
@@
$\mathrm{SLIN}$ [14] $=$ * * 2
$\mathrm{NRE}$ [14] $=$ * * *
@@ 4 Novel results with new ingredients
For virus machines in generating mode, the notation used in previous works [4] was $\mathrm{NVM}(p,q,r)$ where it denotes the family of natural number sets generated by virus machines with at most $p$ hosts, $q$ instructions and $r$ viruses at each instant in the computation. The computational completeness has been proven when these ingredients are unbounded [1], that is, $\mathrm{NRE}=\mathrm{NVM}(*,*,*)$, but the power decreases when one of them is bounded, for example, a characterisation of semilinear sets has been proven in [14], that is, $\mathrm{SLIN}=\mathrm{NVM}(*,*,r)$ for each $r\geq 2$. However, the three ingredients mentioned seem to be poor information. We propose the following notation:
@@
  • $t$ is the maximum $out-degree$ of each host in the host graph,
@@
First, we believe that the instruction-channel graphs have something to say in the sense of computational power, that is why we fix bijection that will show a new frontier of computational power when we combine it with other ingredients for instance in theorems 4.44.9 and  4.11.
Secondly, fixing the directed graphs as trees is a first intention to approach the topology structure of the graphs, fixing them as trees (when $u$ or $v$ are zero) will allow us to fix characterisations to finite sets (such as Corollary 4.3 or Theorem 4.4). Nevertheless, we believe that not only whether or not there is a loop is important, but also the size of the loop. The relation between these sizes and the power of the devices will have interesting results, for instance theorems  4.94.114.12 and 4.18, and Lemma 3.1. Note that $u=1$ is the same as $u=0$ as there are no self-arcs in the host graph.
In addition, the weight of the arcs is our unique way to increase the amount of viruses, fixing this weight to one where the power falls substantially, but if it is $2$ we can get novel universality results (see Theorem 4.18).
@@ 4.1 Finite sets
    The proof is followed by the previous corollary and the inclusions of [4]. In a simple glance, we can see in Figure 2 that both the instruction and the host graphs are trees.
@@
The proof of the first two statements is analogous to the proof of Theorem 4.4. For the third statement, it can be proved by applying the Lemma 3.1 and the Corollary 4.7.
@@ 4.2 Semilinear sets
The authors in [14] characterise semilinear sets by VMs in generating mode, that is, $SLIN=(*,*,r)$, for all $r\geq 2$. Reviewing how the proof was constructed, we can assure the following result with new ingredients:
@@
superset-of-or-equals This part of the proof is equal to the technique applied by the authors in [14]. That is, simulating the right-linear grammar, this is possible due to the amount of viruses of each host at each moment of the computation being bounded, and thus the number of possible configurations is finite. We still are in the same conditions, so the same proof can be made.
@@
[…]
@@ 4.3 Universality
For the proof, we refer to [14], where the Turing completeness was proven by simulating register machines in a modular way. In that simulation, one can see on a simple glance that the weight of the channels are bounded by $2$, thus the theorem is proved.
@@ 4.4 Discussion of new ingredients
Here, table 3 summarises the previous results in addition to the new results obtained in this work with the proposed new ingredients. First, the singleton sets were characterised in the previous section with the old ingredients; note that including the new ingredients is straightforward. With finite sets, it is interesting to note that the new characterisations were previously inclusions, just fixing the host graph and/or the instruction graph to a tree. Showing the strict inclusion with the loops of size two in the host graph and size three in the instruciton graph. These loop restrictions are the same for characterising the semilinear sets, but unbounding the number of hosts instead, note that despite the $\beta$ ingredient seems to be very restrictive, we can characterise this family of sets. The most interesting result of this section arises when we unbound the size of the loops in the instruction graph and $\beta=F$, which allow us to reduce the number of necessary hosts to two. Last but not least, the universality result from [14] has been revised with the new ingredients, showing that there is a huge frontier in computational power, from finite sets with weight $1$, to $\mathrm{NRE}$ with weight $2$.
Family of sets Relation $h$ $i$ $nv$ $wc$ $outd$ $\alpha^{\hbar}$ $\alpha^{\iota}$ $\beta$
Singleton $=$ (Theorem 3.5) 1 1 1 1 1 0 0 T
Finite sets $=$ (Theorem 1) 1 * * 1 1 0 0 F
$=$ (Theorem 2) * * 1 * 1 0 0 T
$=$ (Theorem 4.9) 2 * 2 2 2 2 0 F
@@
$\mathrm{SLIN}$ $=$ (Theorem 4.11) * * 2 * 2 2 3 T
$=$ (Theorem 4.12) 2 * 2 2 2 2 * F
$\mathrm{NRE}$ $=$ (Theorem 4.18) * * * 2 * * * F
@@ 5 Conclusions
Some directions for future work include the extension to parallel VMs e.g. [17, 18]. In such VMs more than one instruction can be active at each computation step, so some of the restrictions in the present work may not apply. Another natural extension of the present work is to apply them to VMs for accepting inputs or as transducers, that is, computing functions using both input and output [4].
The results in normal forms from the present work can better inform the design of applications or implementations of VMs. For instance, perhaps with ideas from [15, 16], sequels of the simulator from [19], or applications of VMs such as edge detection [20], or cryptography [21] can be improved by applying simplifications based on normal forms. Similarly, knowing which types of instruction, host, or channel graphs are used, the reachability of one configuration from another and other properties, can be decidable: in this case, searching for efficient algorithms and applications to formal verification are open problems. The results of the classes below $\mathrm{NRE}$ help inform the computational complexity of problem solving with a model. In the case of VMs, a future direction is to investigate deeper into “sub-$NRE$” classes: in this way a better view of efficiency or lack of it, can be given for VMs. Lastly, new and optimal lower bounds are expected to improve the results of the present work, summarised in Table 3. New normal forms can perhaps include new ingredients or semantics not previously considered, such as a deeper focus on deterministic versus nondeterministic computations.

### v3

@@ 3.5 Discussion of old ingredients
Singleton (Theorem 3.5) == 1 1 1
$\mathrm{NLinFIN}$ (Proposition 3.7) == 2 2 *
$\mathrm{NCombFIN}$ (Proposition 3.9) == 3 3 *
Finite sets [GCRVM] $\subseteq$ 1 * *
[GCRVM] $\subseteq$ * * 1
@@
$\mathrm{SLIN}$ [computing-VM] == * * 2
$\mathrm{NRE}$ [computing-VM] == * * *
@@ 4 Novel results with new ingredients
For virus machines in generating mode, the notation used in previous works [GCRVM] was $\mathrm{NVM}(p,q,r)$ where it denotes the family of natural number sets generated by virus machines with at most $p$ hosts, $q$ instructions and $r$ viruses at each instant in the computation. The computational completeness has been proven when these ingredients are unbounded [VM], that is, $\mathrm{NRE}=\mathrm{NVM}(*,*,*)$, but the power decreases when one of them is bounded, for example, a characterisation of semilinear sets has been proven in [computing-VM], that is, $\mathrm{SLIN}=\mathrm{NVM}(*,*,r)$ for each $r\geq 2$. However, the three ingredients mentioned seem to provide poor information about the computation. Thus, we propose the following notation:
@@
  • $t$ is the maximum outdegreeout-degree of each host in the host graph,
@@
First, we believe that the instruction-channel graphs have something to say in the sense of computational power, that is why we fix the bijection that will show a new frontier of computational power when we combine it with other ingredients, for instance, in theorems 4.44.9 and  4.11.
Secondly, fixing the directed graphs as trees (when $u$ or $v$ are zero) is a first intention to approach the topology structure of the graphs, as they allow us characterisations of finite sets (such as Corollary 4.3 or Theorem 4.4). Nevertheless, we believe that not only whether or not there is a loop is important, but also the size of the loop. The relation between these sizes and the power of the devices will have interesting results, for instance, Theorems  4.94.114.12, and 4.18, and Lemma 3.1. Note that $u=1$ is the same as $u=0$ as there are no self-arcs in the host graph.
Thirdly, the weight of the arcs is the unique way to increase the amount of viruses, fixing this weight to one makes the power falls substantially (only finite sets as seen in Theorem 1), but if it is $2$ we can get novel universality results (see Theorem 4.18).
@@ 4.1 Finite sets
    The proof is followed by the previous corollary and the inclusions of [GCRVM]. In a simple glance, we can see in Figure 2 that both the instruction and the host graphs are trees.
@@
The proof of the first two statements is analogous to the proof of Theorem $4.$ For the third statement, it can be proved by applying the Lemma 3.1 and the Corollary 4.7.
@@ 4.2 Semilinear sets
The authors in [computing-VM] characterise semilinear sets by VMs in generating mode, that is, SLIN=(,,r)SLIN=(*,*,r), for all $r\geq 2$. Reviewing how the proof was constructed, we can assure the following result with new ingredients:
@@
This part of the proof is equal to the technique applied by the authors in [computing-VM]. That is, simulating the right-linear grammar, this is possible due to the amount of viruses of each host at each moment of the computation being bounded, and thus the number of possible configurations is finite. We still are in the same conditions, so the same proof can be made.
@@
Q1Q2As both $Q_{1},Q_{2}$ are arithmetic progression, we can apply the Lemma 4.14, we will denote that $Q_{j}$ is generated by $\Pi_{j}$ for each $j\in\{1,2\}$, this sub-index notation is extended to the heterogeneous networks of each machine. Note that the host graphs and the initial amount of viruses of each virus machine remains equal. The construction is visually explained in Figure 13. From which it can be seen in a simple glance that it generates $Q_{1}\cup Q_{2}$, proving that $Q_{1}\cup Q_{2}$ is in the families of the statement.
j=1m1QjQmFor the inductive step, let us suppose that a virus machine $\Pi_{m-1}$ generates the union with host graph mentioned in the base case. Applying again Lemma 4.14, let $\Pi_{m}$ the virus machine that generates the set $Q_{m}$. The construction of the virus machine that generates the union of both sets will be analogous to the previous one, visually explained in Figure 14, thus it has been proved that $\cup_{j=1}^{m}Q_{j}$ is in the families of the statement.
@@ 4.3 Universality
For the proof, we refer to [computing-VM], where the Turing completeness was proven by simulating register machines in a modular way. In that simulation, one can see on a simple glance that the weight of the channels are bounded by $2$, thus the theorem is proved.
@@ 4.4 Discussion of new ingredients
Here, table 3 summarises the previous results in addition to the new results obtained in this work with the proposed new ingredients. First, the singleton sets were characterised in the previous section with the old ingredients; note that including the new ingredients is straightforward. With finite sets, it is interesting to note that the new characterisations were previously inclusions, just fixing the host graph and/or the instruction graph to a tree. Showing the strict inclusion with the loops of size two in the host graph and size three in the instruciton graph. These loop restrictions are the same for characterising the semilinear sets, but unbounding the number of hosts instead. Note that despite the $\beta$ ingredient seems to be very restrictive, we can characterise this family of sets. The most interesting result of this section arises when we unbound the size of the loops in the instruction graph and $\beta=F$, which decreases the number of necessary hosts to two. Last but not least, the universality result from [computing-VM] has been revised with the new ingredients, showing that there is a huge frontier in computational power, from finite sets with weight $1$, to $\mathrm{NRE}$ with weight $2$.
Family of sets Relation $h$ $i$ nvnv wcwc outdoutd $\alpha^{\hbar}$ $\alpha^{\iota}$ $\beta$
Singleton == (Theorem 3.5) 1 1 1 1 1 0 0 T
Finite sets == (Theorem 1) 1 * * 1 1 0 0 F
== (Theorem 2) * * 1 * 1 0 0 T
== (Theorem 4.9) 2 * 2 2 2 2 0 F
@@
$\mathrm{SLIN}$ == (Theorem 4.11) * * 2 * 2 2 3 T
== (Theorem 4.12) 2 * 2 2 2 2 * F
$\mathrm{NRE}$ == (Theorem 4.18) * * * 2 * * * F
@@ 5 Conclusions
Some directions for future work include the extension to parallel VMs e.g. [vm_workflow2024, Antonio2024a]. In such VMs more than one instruction can be active at each computation step, so some of the restrictions in the present work may not apply. Another natural extension of the present work is to apply them to VMs for accepting inputs or as transducers, that is, computing functions using both input and output [GCRVM].
The results in normal forms from the present work can better inform the design of applications or implementations of VMs. For instance, perhaps with ideas from [vm_matriz_bwmc2024, vm_matriz2024], sequels of the simulator from [vm_sim], or applications of VMs such as edge detection [Antonio2024b], or cryptography [CryptoVM] can be improved by applying simplifications based on normal forms. Similarly, knowing which types of instruction, host, or channel graphs are used, the reachability of one configuration from another and other properties, can be decidable: in this case, searching for efficient algorithms and applications to formal verification are open problems. The results of the classes below $NRE$ help inform the computational complexity of problem solving with a model. In the case of VMs, a future direction is to investigate deeper into “sub-NRENRE” classes: in this way a better view of efficiency or lack of it, can be given for VMs. Lastly, new and optimal lower bounds are expected to improve the results of the present work, summarised in Table 3. New normal forms can perhaps include new ingredients or semantics not previously considered. For instance, a deeper focus on deterministic versus nondeterministic computations, as well classes and trade-offs for time and space.